SYM:

7a-4=18 im not usually doing this idk what to do help? normally it's a lot easier

jkmiec26:

Ok so you'll want to isolate the a. You'll start by adding 4 to both sides. Giving you 7a=22. Then you'll want to divide 7 by 22 and you'll get a decimal. Which is 3.14.... etc. 

SYM:

@jkmiec26 wrote:
Ok so you'll want to isolate the a. You'll start by adding 4 to both sides. Giving you 7a=22. Then you'll want to divide 7 by 22 and you'll get a decimal. Which is 3.14.... etc. Created with Raphaƫl7a-4=18+4+47a=2277a=3.14...Reply Using Drawing
thank you so much I think I get it now

SYM:

sry i was late my internet is crap